OpenStudy (mgraham3641):

-4x-2y=20 2x+y=19 this is a systems using elimination

OpenStudy (shadowlegendx):

Do you know how to do elimination?

OpenStudy (mgraham3641):

kind of we have been learning it for only like 2 days

OpenStudy (shadowlegendx):

Have you guys gone over one solution, infinite solutions, and no solutions?

OpenStudy (mgraham3641):

yes

OpenStudy (shadowlegendx):

Good, I'll show you why I asked that

OpenStudy (shadowlegendx):

First step to elimination, is cancelling out one variable. The first thing I see is -4x and 2x, which is perfect. Our negative is already lined up. \[-4x - 2y = 20\] \[2x + y = 19\] Multiply the second equation by 2 \[-4x - 2y = 20\] \[4x + 2y = 38\] Add the two equations Both x and y cancel \[0 = 58\]

OpenStudy (mgraham3641):

give me a little bit to figure this out

OpenStudy (shadowlegendx):

What part are you confused on?

OpenStudy (mgraham3641):

im not really confused its just a lot of numbers

OpenStudy (mgraham3641):

and we really never learned to multiply the equations ?

OpenStudy (shadowlegendx):

Do you understand why we're allowed to multiply the equations?